The market for lemon has 10potential consumers ,each having an individual demand curve ,p =101-10Q,where p is price in a dollars per cup and Qi is the number of cups demanded per week by the ith consumer . Find the market demand curve using algebra.Draw an individual demand curve and the market demand curve.What is the quantity demand by each consumer and in the market as a whole when lemon is priced at p=$1/cup?


1
Expert's answer
2021-02-03T16:09:52-0500

"Q = \sum_{i=1}^{10}q_i \\

= q_1+q_2+q_3+...+q_{10} \\

= (\frac{101-P}{10})_1 + (\frac{101-P}{10})_2 + (\frac{101-P}{10})_3 + … + (\frac{101-P}{10})_{10} \\

= (10.1 -0.1P)_1 + (10.1 -0.1P)_2 + (10.1 -0.1P)_3 + … + (10.1 -0.1P)_{10} \\

= 10 \times (10.1 -0.1P) \\

= 101 -P \\

P = 101 - Q"

This is the market demand curve for lemonade.

Given the individual demand curve and market demand curve, the graphs will be as follows:



The market demand curve is flatter than the individual demand curve.

Given the price of one cup of lemonade, the individual demand is calculated as follows:

"P = 101 - 10Q_i \\

Q_i = \frac{101-1}{10} = 10"

Hence, the individual demand for lemonade is 10 cups at $1.

The market demand is calculated as follows:

"P = 101 -Q \\

Q = 101 -1 = 100"

Hence, the market demand for lemonade is 100 cups at $1.